# Contact Wasserstein Geodesics for Non-Conservative Schrödinger Bridges

### v2

@@ D.4 Time Complexity and Practical Considerations
to a batch of samples. As discussed in Sec. 3, the Fisher information $I({\rho}^{t_{k}}_{\theta})$ in the potential energy originates from the entropy regularization in the SB formulation and does not require explicit computation. Its effect is implicitly captured by the entropy-regularized Wasserstein distance. Therefore, evaluating $\Phi^{t_{k}}$ reduces to computing the scalar functions $H$ and $U$, along with estimating the entropy term $\log(2{\rho}^{t_{k}}_{\theta})\,{\rho}^{t_{k}}_{\theta}$, which is the computational bottleneck. This term can be estimated using a $k$-NN entropy estimator with time complexity $\mathcal{O}(D\,N\log N)$ (biophysica2040031). Considering all components, and given that $K\geq M$, the overall time complexity becomes,
@@
Our method demonstrates highly favorable scaling properties, offering a significant advantage over existing approaches. Notably, its computational complexity scales linearly with data dimensionality $D$ and nearly linearly with the batch size $N$. This stands in stark contrast to existing methods like (hong2025trajectory), which scale quadratically in $N$ and potentially exponentially in $D$. Furthermore, our model’s performance is only weakly influenced by the number of marginals and it circumvents the expensive outer iteration loops required for convergence in methods like (chen2023deep; shen2025multimarginal).
Appendix E Implementation Details
E.1 Empirical Approximation of the Wasserstein-2 Distance
@@
represents the squared Euclidean distance between sample $x_{c,i}$ and sample $x_{d,j}$. A transport plan is then defined as a matrix $\pi\in\mathbb{R}_{+}^{N\times M}$ that assigns how much mass to move from each $x_{c,i}$ to each $x_{d,j}$, minimizing the total transport cost weighted by $C$. To avoid degenerate solutions where all mass is concentrated on a few points, entropy regularization is introduced, encouraging smoother and more distributed transport plans. For this computation, we employ the SamplesLoss function from the geomloss library (feydy2020geometric), with parameters resumed in Table 9.
@@ E.2 Teaser Image
  • The concentration of the discretization nodes in regions of higher energy (i.e., where $H(\rho^{t_{k}}_{\theta})$ is larger). As the evolution of the Hamiltonian can be prescribed, selecting a decreasing or increasing law results in a stochastic trajectory that slows down or speeds up over time, respectively, as discussed in Appendix D.3.
@@
Appendix F Extended Results
F.1 Experimental Setup
The LiDAR Manifold Navigation and Cell Sequencing experiments were conducted on a machine equipped with 13th Gen Intel® Core i7-13850HX CPUs. The Image Generation experiment was run on a system with an NVIDIA GeForce RTX5090 GPU (32GB VRAM, CUDA12.9, driver version 575.64.03). Results for the LiDAR Manifold Navigation (Table 6), Single Cell Sequencing (Table 6), Sea Temperature Prediction (Table 6 and Appendix F.4), Robot Task Reconstruction (Table 6), and Unpaired Image Transfer (Table 7) are based on a total of ten evaluations obtained from training runs with different initial conditions. The tables report the mean and standard deviation of these distributions.
@@ F.1 Experimental Setup
Parameter Lidar Cell Sea Robot Transfer
# samples NN $1000$ $1000$ $200$ $100$ $1000$
Weight $w_{b}$ $10$ $10$ $100$ $10$ 11
Weight $w_{m}$ - $10$ $100$ - -
Weight $w_{g}$ 11 11 11 11 11
@@
Lidar Cell Transfer
@@
Sea Robot
Number of blocks $K$ 55 88
Input channels 11 33
Output channels 11 33
Layers per block 11 22
@@
Step size $\tau$ 11 11
Input dimension $d$ $4096$ $12288$
@@
Experiment DSBM GSBM SBIRR DM-SB CWG
LiDAR Navigation 0.14M0.14\ $\text${M} 0.33M0.33\ $\text${M} - - 0.017M0.017\ $\text${M}
Cell Sequencing - - 0.013M0.013\ $\text${M} 2.5M2.5\ $\text${M} 0.017M0.017\ $\text${M}
Sea Temperature 79.2M79.2\ $\text${M} 271M271\ $\text${M} - - 5.2M5.2\ $\text${M}
Robot Task 79.2M79.2\ $\text${M} 271M271\ $\text${M} - - 7.3M7.3\ $\text${M}
Unpaired Transfer 5.7M5.7\ $\text${M} 5.7M5.7\ $\text${M} - - 15.7M15.7\ $\text${M}
@@
F.2 LiDAR Manifold Navigation
The LiDAR dataset (opentopo_rainier_2025) consists of point clouds contained in the domain $5,5]^{3}\subset\mathbb{R}^{3}$. The objective of the experiment is to construct a bridge across the data manifold for connecting two distributions while avoiding regions of high elevation and remaining closely aligned with the manifold structure. The initial distribution ${\rho}_{a}$ is composed by a mixture of $4$ Gaussian distributions, the target distribution ${\rho}_{b}$ is composed of $2$ Gaussians on the two sides of the mountain. The manifold shape is incorporated in the problem through the potential function $U$, inherited from the baseline (liu2024generalized),
@@ F.2 LiDAR Manifold Navigation
Here, $\psi(x)$ denotes the projection of a point $x$ onto an approximate tangent plane, estimated from its $p$ nearest neighbors on the data manifold. The notation $\psi^{(z)}(x)$ refers to the $z$-coordinate of $\psi(x)$, i.e., the height of the fitted plane. The weights $w_{\text{manifold}}$ and $w_{\text{height}}$ control the relative importance of the two terms in the potential function. We now detail the construction of $\psi(x)$. Let $N_{p}(x)=\{x^{l}_{1},\dots,x^{l}_{p}\}$ denote the set of $p$ nearest neighbors of $x\in\mathbb{R}^{3}$ in the dataset. To approximate the local tangent plane, we employ a moving least-squares (MLS) procedure (levin1998approximation). Specifically, the plane parameters $(a,b,c)$ are obtained by solving,
@@
The quantitative results reported in Table 6 aim to characterize the two key aspects of the Schrödinger Bridge problem, as formulated in the stochastic optimization problems (2) and (4). The Optimality metric reports the length of the stochastic bridge, computed as the square root of the cost functional $J$. It quantifies the ability of the approach to find low transport-cost solutions connecting the marginals $\rho_{a}$ and $\rho_{b}$. In contrast, the Feasibility metric is computed as the Wasserstein distance between the marginals and the nodes of the discretized bridge that approximate them. This metric measures how well the boundary constraints in the optimization problems are satisfied. Formally, these metrics are defined as,
@@
F.3 Single Cell Sequencing
The Embryoid Body (EB) stem cell differentiation dataset (moon2019visualizing) captures cell state progression across five developmental stages $t_{0},t_{1},t_{2},t_{3},t_{4}]$ over a 27-day period. Snapshots were collected at five discrete time intervals: $t_{0}\in[0,3]$, $t_{1}\in[6,9]$, $t_{2}\in[12,15]$, $t_{3}\in[18,21]$, and $t_{4}\in[24,27]$. These stages involve significant structural changes, with cells moving and reorganizing within increasingly stiff tissue while consuming and releasing mechanical energy (zeevaert2020cell; kinney2014mesenchymal). Consequently, the resulting dynamics exhibit energy dissipation and are better described by the NCGSB framework than by energy-conserving models. In this experiment, we evaluate the framework’s ability to generalize to regions with no available data by dividing the dataset into a training set $t_{0},t_{2},t_{4}]$ and a validation set $t_{1},t_{3}]$. The geometry of the data manifold is incorporated into the NCGSB problem through a potential function $U$, defined as,
@@ F.3 Single Cell Sequencing
where $x^{t}_{j}\sim{\rho}^{t}$ denotes a sample from the posterior distribution, and $x_{i}\sim\{{\rho}_{a},{\rho}_{m},{\rho}_{b}\}$ are samples from the marginal distributions. $N_{1}$ and $N_{2}$ indicate the number of samples taken from ${\rho}^{t}$ and $\{{\rho}_{a},{\rho}_{m},{\rho}_{b}\}$, respectively, while $\gamma$ represents a spatial scaling parameter. The exponential term acts as a kernel-like weight that reduces the influence in the inverse summation $\sum_{i=1}^{N_{1}}[\cdot]^{-1}$ of the data points far from the bridge. Globally, the potential function $U$ measures the distance of the bridge ${\rho}^{t}$ from the available data $\{{\rho}_{a},{\rho}_{m},{\rho}_{b}\}$ (the training set) and its minimization guides the construction of a bridge that stays close to the known manifold while generalizing effectively to regions without data (the validation set). While other approaches leveraging this dataset (tong2020trajectorynet; shen2025multimarginal) first embed the data into a $100$-dimensional feature space using principal component analysis (PCA) and then restrict the analysis to the first five dimensions, this procedure excessively linearizes and flattens the data manifold, making navigation trivial and eliminating the need for intermediate marginals (shen2025multimarginal). To better preserve the manifold’s geometry, we instead apply the PHATE algorithm (moon2019visualizing) to the $100$-dimensional representation, producing a $5$-dimensional nonlinear embedding that more faithfully captures the original structure.
In the extensive results reported in Tables 18 and 18, corresponding to the cell dynamics reconstruction shown in Fig. 12, the scalar Hamiltonian function $H^{t_{k}}$ in the pullback Jacobi metric, $T_{\theta}^{*}\tilde{g}_{\mathrm{J}}=\big(H^{t_{k}}-\mathcal{F}(\rho^{t})-\mathcal{B}(\rho^{t})\big)\,T_{\theta}^{*}g^{\mathcal{W}_{2}}$ used in the CWG approach, is linearly varied from an initial value $H^{t_{0}}=0.82$ to a final value $H^{t_{K}}=1$, along the trajectory defined by the first principal dimension $D_{1}$, i.e.,
@@
This law is treated as a hyperparameter of the methodology. All ten experiments reported are conducted under this Hamiltonian behavior. This increasing trend, associated with the reduction of physical energy (see Appendix D.3), aligns with the underlying biological process and has been shown to produce better results than the energy-conserving case, where $H^{t_{k}}=1$ is kept constant (see Table 20 and Figure 13(a)). As detailed in Appendix E.2, the behavior of $H^{t_{k}}$ plays a crucial role in determining how quickly the nodes of the stochastic path depart from the initial reference distribution and approach the subsequent reference marginal. In Figure 13(b), the dataset exhibits a decreasing Wasserstein distance $d_{\mathcal{W}_{2}}$ between successive cell snapshots, reflecting that cell differentiation progresses more rapidly at early stages (zeevaert2020cell). Training with the Hamiltonian behavior defined in equation (64) enables the model to capture this trend more accurately.
@@ F.4 Sea Temperature Prediction
The NOAA OISST v2 High Resolution Dataset (huang2021improvements) is a long-term Climate Data Record that integrates observations from multiple platforms (satellites, ships, buoys, and Argo floats) into a global gridded product. For this experiment, we use daily averages of sea surface temperature in the Gulf of Mexico between $1981$ and $2024$, represented as $64\times 64$ single-channel images. We cluster the measurements over five-year periods and select five representative months to define five time frames: January ($t_{0}$), March ($t_{1}$), May ($t_{2}$), July ($t_{3}$), and September ($t_{4}$). Each month corresponds to a distribution of images, denoted as $\{{\rho}_{a}\ \text{(January)},{\rho}_{m_{1}}\ \text{(March)},{\rho}_{m_{2}}\ \text{(May)},{\rho}_{m_{3}}\ \text{(July)},{\rho}_{b}\ \text{(September)}\}$. A sample from one of these distributions is a heatmap of the Gulf’s temperature for a specific day in the corresponding month of the specified five-year period. The goal of this test is to evaluate our method’s ability to interpolate across missing time frames, generating realistic temperature maps for months without data. To this end, we partition the dataset into a training set $\{t_{0},t_{2},t_{4}\}$ and a validation set $\{t_{1},t_{3}\}$, and assess the quality of predictions on the held-out months. To encourage generalization beyond the training data, we introduce a potential function $U$ that penalizes deviations from the learned data manifold. Building on the approach of song2025riemanniangeometric, where generative models are evaluated by measuring the distance between their outputs and a geometric manifold of real images learned by a VAE, we adopt a similar strategy. Specifically, we use a state-of-the-art VAE architecture, with parameters listed in Table 21, to learn the manifold of the training images in our dataset. The potential function $U(x^{t})$, for samples $x^{t}\sim{\rho}^{t}$, is then defined as the squared distance between a bridge sample $x^{t}$ and its VAE-projected reconstruction $\tilde{x}^{t}=\text{VAE}(x^{t})$: $U(x^{t})=\lVert x^{t}-\tilde{x}^{t}\rVert^{2}$. Results for each five-year periods are presented below.
@@
In these results, the scalar Hamiltonian function $H^{t_{k}}$ in the pullback Jacobi metric, $T_{\theta}^{*}\tilde{g}_{\mathrm{J}}=\big(H^{t_{k}}-\mathcal{F}(\rho^{t})-\mathcal{B}(\rho^{t})\big)\,T_{\theta}^{*}g^{\mathcal{W}_{2}}$ is linearly varied from an initial value $H^{t_{0}}=1.36$ to a final value $H^{t_{K}}=1.$. This law is treated as a hyperparameter of the methodology, and all ten experiments reported in the following tables were conducted under this Hamiltonian behavior. In the geodesic computation, the solution path seeks to minimize $T_{\theta}^{*}\tilde{g}_{\mathrm{J}}$. The functional $\mathcal{F}(\rho^{t})$ measures the distance with respect to the data manifold, i.e., it evaluates whether the samples from the bridge remain coherent with the images observed during training, and is minimized along the trajectory. Since $H^{t_{k}}$ is prescribed, the entropy term $\mathcal{B}(\rho^{t})$ is effectively maximized up to the limit set by $H^{t_{k}}$, as the metric cannot become negative. Hence, the assigned Hamiltonian energy $H^{t_{k}}$ determines the admissible entropy level of the bridge. In this experiment, a decreasing trend of $H^{t_{K}}$, corresponding to an increase in physical energy (see Appendix D.3), was observed to be beneficial for modeling the warmer months, which appear to be more distant from the rest of the dataset in terms of Wasserstein distances compared to the colder months. A higher final energy $H^{t_{K}}$, and thus a higher final entropy $\mathcal{B}$, was observed to be beneficial for modeling the warmer months. This effect can be associated with the increased thermodynamic entropy of such cases, leading to more diverse samples and larger variability across the data manifold. By prescribing a higher energy level, the model is encouraged to capture this diversification, exploring regions of the data manifold that in other contexts arvanitidis2018latent are regarded as uncertain and are typically avoided.
@@
Table 21: Architecture of the ConvVAE used in the Sea Temperature Prediction experiment. All Conv2D and ConvTranspose2D layers use kernel size 44, stride 22, and padding 11, followed by ReLU activations (except the last layer, which uses Sigmoid).
@@
Metric CWG GSBM DSBM
@@
Metric CWG GSBM DSBM
@@
Metric CWG GSBM DSBM
@@
Metric CWG GSBM DSBM
@@
Metric CWG GSBM DSBM
@@
Metric CWG GSBM DSBM
@@
Metric CWG GSBM DSBM
@@
Metric CWG GSBM DSBM
@@
Metric CWG GSBM DSBM

### v3

@@ E.5 Time Complexity and Practical Considerations
to a batch of samples. As discussed in Sec. 3, the Fisher information $I({\rho}^{t_{k}}_{\theta})$ in the potential energy originates from the entropy regularization in the SB formulation and does not require explicit computation. Its effect is implicitly captured by the entropy-regularized Wasserstein distance. Therefore, evaluating $\Phi^{t_{k}}$ reduces to computing the scalar functions $H$ and $U$, along with estimating the entropy term $\log(2{\rho}^{t_{k}}_{\theta})\,{\rho}^{t_{k}}_{\theta}$, which is the computational bottleneck. This term can be estimated using a $k$-NN entropy estimator with time complexity $\mathcal{O}(D\,N\log N)$ (Borelli et al., 2022). Considering all components, and given that $K\geq M$, the overall time complexity becomes,
@@
Our method demonstrates highly favorable scaling properties, offering a significant advantage over existing approaches. Notably, its computational complexity scales linearly with data dimensionality $D$ and nearly linearly with the batch size $N$. Furthermore, our model’s performance is only weakly influenced by the number of marginals and it circumvents the expensive outer iteration loops. This stands in stark contrast to existing methods, as elaborated next:
  • Hong et al. (2025) scales quadratically in $N$ and exponentially in $D$, with time complexity,
    $\mathcal{O}\!\left(T_{\text{iter}}\,DKN^{2}\exp\!\bigl(\alpha(D+1)\bigr)\right),$ (56)
    where $T_{\text{iter}}$ denotes the number of Belief Propagation iterations required for convergence, and $\alpha$ is a parameter determined by the order of the Gaussian process and structural assumptions on their representation basis.
  • Chen et al. (2023) proposed a method with time complexity,
    $\mathcal{O}\!\left(T_{\text{iter}}KN(2D)^{2}\right),$ (57)
    where $T_{\text{iter}}$ denotes the number of Bregman iterations needed for convergence. The quadratic dependence on $2D$ arises from the state space being doubled by incorporating velocity, which comes at the cost of computational efficiency.
  • Shen et al. (2025) employed a score-matching SB solver (Vargas et al., 2021) for the mmSB problem with computational complexity,
    $\mathcal{O}\!\left(T_{\text{iter}}\,MN(2D)\right),$ (58)
    where $T_{\text{iter}}$ again denotes the number of Bregman iterations needed for convergence, with reported values between $2000$ and $4000$. Although the method is nearly linear in both $N$ and $D$, it depends explicitly on the number of intermediate marginals $M$ and suffers from additional outer-loop iterations.
As highlighted in this analysis, our CWG method advances the state of the art as a remarkably efficient computational approach: it scales linearly in all relevant terms and does not require iterations. The baselines (Shi et al., 2023; Bortoli et al., 2024; Liu et al., 2024) do not explicitly report their computational efficiency, so the comparison with our method is addressed empirically in the experiments.
Appendix F Implementation Details
F.1 Empirical Approximation of the Wasserstein-2 Distance
@@
represents the squared Euclidean distance between sample $x_{c,i}$ and sample $x_{d,j}$. A transport plan is then defined as a matrix $\pi\in\mathbb{R}_{+}^{N\times M}$ that assigns how much mass to move from each $x_{c,i}$ to each $x_{d,j}$, minimizing the total transport cost weighted by $C$. To avoid degenerate solutions where all mass is concentrated on a few points, entropy regularization is introduced, encouraging smoother and more distributed transport plans. For this computation, we employ the SamplesLoss function from the geomloss library (Feydy, 2020), with parameters resumed in Table 10.
@@ F.2 Teaser Image
  • The concentration of the discretization nodes in regions of higher energy (i.e., where $H(\rho^{t_{k}}_{\theta})$ is larger). As the evolution of the Hamiltonian can be prescribed, selecting a decreasing or increasing law results in a stochastic trajectory that slows down or speeds up over time, respectively, as discussed in Appendix E.4.
@@
Appendix G Extended Results
G.1 Experimental Setup
The LiDAR Manifold Navigation and Cell Sequencing experiments were conducted on a machine equipped with 13th Gen Intel® Core i7-13850HX CPUs. The Image Generation experiment was run on a system with an NVIDIA GeForce RTX5090 GPU (32GB VRAM, CUDA12.9, driver version 575.64.03). Results for the LiDAR Manifold Navigation (Table 3), Single Cell Sequencing (Table 3), Sea Temperature Prediction (Table 4 and Appendix G.4), Robot Task Reconstruction (Table 6), FFHQ Transfer (Table 7), and MNIST-to-EMNIST (Table 38) experiments are based on a total of ten evaluations obtained from training runs with different initial conditions. The tables report the mean and standard deviation of these distributions.
@@ G.1 Experimental Setup
Parameter Lidar Cell Sea Robot FFHQ Mnist
# samples NN $1000$ $1000$ $200$ $100$ $1000$ $300$
Weight $w_{b}$ $10$ $10$ $100$ $10$ 11 11
Weight $w_{m}$ - $10$ $100$ - - -
Weight $w_{g}$ 11 11 11 11 11 11
@@
Lidar Cell FFHQ
@@
Sea Robot Mnist
Number of blocks $K$ 55 88 66
Input channels 11 33 11
Output channels 11 33 11
Layers per block 11 22 22
@@
Step size $\tau$ 11 11 11
Input dimension $d$ $4096$ $12288$ $784$
@@
Experiment DSBM SB-Flow GSBM SBIRR DM-SB CWG
LiDAR Navigation 0.14M0.14\ $\text${M} - 0.33M0.33\ $\text${M} - - 0.017M0.017\ $\text${M}
Cell Sequencing - - - 0.013M0.013\ $\text${M} 2.5M2.5\ $\text${M} 0.017M0.017\ $\text${M}
Sea Temperature 79.2M79.2\ $\text${M} 11M11\ $\text${M} 271M271\ $\text${M} - - 5.2M5.2\ $\text${M}
Robot Task 79.2M79.2\ $\text${M} 11M11\ $\text${M} 271M271\ $\text${M} - - 7.3M7.3\ $\text${M}
Unpaired Transfer 5.7M5.7\ $\text${M} 6.6M6.6\ $\text${M} 5.7M5.7\ $\text${M} - - 15.7M15.7\ $\text${M}
Mnist-to-Emnist 4.6M4.6\ $\text${M} 3.3M3.3\ $\text${M} 4.6M4.6\ $\text${M} - - 5.2M5.2\ $\text${M}
@@
G.2 LiDAR Manifold Navigation
The LiDAR dataset (OpenTopography, 2025) consists of point clouds contained in the domain $5,5]^{3}\subset\mathbb{R}^{3}$. The objective of the experiment is to construct a bridge across the data manifold for connecting two distributions while avoiding regions of high elevation and remaining closely aligned with the manifold structure. The initial distribution ${\rho}_{a}$ is composed by a mixture of $4$ Gaussian distributions, the target distribution ${\rho}_{b}$ is composed of $2$ Gaussians on the two sides of the mountain. The manifold shape is incorporated in the problem through the potential function $U$, inherited from the baseline (Liu et al., 2024),
@@ G.2 LiDAR Manifold Navigation
Here, $\psi(x)$ denotes the projection of a point $x$ onto an approximate tangent plane, estimated from its $p$ nearest neighbors on the data manifold. The notation $\psi^{(z)}(x)$ refers to the $z$-coordinate of $\psi(x)$, i.e., the height of the fitted plane. The weights $w_{\text{manifold}}$ and $w_{\text{height}}$ control the relative importance of the two terms in the potential function. We now detail the construction of $\psi(x)$. Let $N_{p}(x)=\{x^{l}_{1},\dots,x^{l}_{p}\}$ denote the set of $p$ nearest neighbors of $x\in\mathbb{R}^{3}$ in the dataset. To approximate the local tangent plane, we employ a moving least-squares (MLS) procedure (Levin, 1998). Specifically, the plane parameters $(a,b,c)$ are obtained by solving,
@@
The quantitative results reported in Table 3 aim to characterize the two key aspects of the Schrödinger Bridge problem, as formulated in the stochastic optimization problems (2) and (4). The Optimality metric reports the length of the stochastic bridge, computed as the square root of the cost functional $J$. It quantifies the ability of the approach to find low transport-cost solutions connecting the marginals $\rho_{a}$ and $\rho_{b}$. In contrast, the Feasibility metric is computed as the Wasserstein distance between the marginals and the nodes of the discretized bridge approximating them. This metric measures how well the boundary constraints in the optimization problems are satisfied. Formally, these metrics are defined as,
@@
G.3 Single Cell Sequencing
The Embryoid Body (EB) stem cell differentiation dataset (Moon et al., 2019) captures cell state progression across five developmental stages $t_{0},t_{1},t_{2},t_{3},t_{4}]$ over a 27-day period. Snapshots were collected at five discrete time intervals: $t_{0}\in[0,3]$, $t_{1}\in[6,9]$, $t_{2}\in[12,15]$, $t_{3}\in[18,21]$, and $t_{4}\in[24,27]$. These stages involve significant structural changes, with cells moving and reorganizing within increasingly stiff tissue while consuming and releasing mechanical energy (Zeevaert et al., 2020; Kinney et al., 2014). Consequently, the resulting dynamics exhibit energy dissipation and are better described by the NCGSB framework than by energy-conserving models. In this experiment, we evaluate the framework’s ability to generalize to regions with no available data by dividing the dataset into a training set $t_{0},t_{2},t_{4}]$ and a validation set $t_{1},t_{3}]$. The geometry of the data manifold is incorporated into the NCGSB problem through a potential function $U$, defined as,
@@ G.3 Single Cell Sequencing
where $x^{t}_{j}\sim{\rho}^{t}$ denotes a sample from the posterior distribution, and $x_{i}\sim\{{\rho}_{a},{\rho}_{m},{\rho}_{b}\}$ are samples from the marginal distributions. $N_{1}$ and $N_{2}$ indicate the number of samples taken from ${\rho}^{t}$ and $\{{\rho}_{a},{\rho}_{m},{\rho}_{b}\}$, respectively, while $\gamma$ represents a spatial scaling parameter. The exponential term acts as a kernel-like weight that reduces the influence in the inverse summation $\sum_{i=1}^{N_{1}}[\cdot]^{-1}$ of the data points far from the bridge. Globally, the potential function $U$ measures the distance of the bridge ${\rho}^{t}$ from the available data $\{{\rho}_{a},{\rho}_{m},{\rho}_{b}\}$ (the training set) and its minimization guides the construction of a bridge that stays close to the known manifold while generalizing effectively to regions without data (the validation set). While other approaches leveraging this dataset (Tong et al., 2020; Shen et al., 2025) first embed the data into a $100$-dimensional feature space using principal component analysis (PCA) and then restrict the analysis to the first five dimensions, this procedure excessively linearizes and flattens the data manifold, making navigation trivial and eliminating the need for intermediate marginals (Shen et al., 2025). To better preserve the manifold’s geometry, we instead apply the PHATE algorithm (Moon et al., 2019) to the $100$-dimensional representation, producing a $5$-dimensional nonlinear embedding that more faithfully captures the original structure.
In the results reported in Tables 19 and 19, corresponding to the cell dynamics reconstruction shown in Fig. 12, the scalar Hamiltonian function $H^{t_{k}}$ in the pullback Jacobi metric, $T_{\theta}^{*}\tilde{g}_{\mathrm{J}}=\big(H^{t_{k}}-\mathcal{F}(\rho^{t})-\mathcal{B}(\rho^{t})\big)\,T_{\theta}^{*}g^{\mathcal{W}_{2}}$ used in the CWG approach, is linearly varied from an initial value $H^{t_{0}}=0.82$ to a final value $H^{t_{K}}=1$, along the trajectory defined by the first principal dimension $D_{1}$, i.e.,
@@
This law is treated as a hyperparameter of the methodology. All ten experiments reported are conducted under this Hamiltonian behavior. This increasing trend, associated with the reduction of physical energy (see Appendix E.4), aligns with the underlying biological process and has been shown to produce better results than the energy-conserving case, where $H^{t_{k}}=1$ is kept constant (see Table 21 and Figure 13(a)). As detailed in Appendix F.2, the behavior of $H^{t_{k}}$ plays a crucial role in determining how quickly the nodes of the stochastic path depart from the initial reference distribution and approach the subsequent reference marginal. In Figure 13(b), the dataset exhibits a decreasing Wasserstein distance $d_{\mathcal{W}_{2}}$ between successive cell snapshots, reflecting that cell differentiation progresses more rapidly at early stages (Zeevaert et al., 2020). Training with the Hamiltonian behavior defined in equation (70) enables the model to capture this trend more accurately.
@@ G.4 Sea Temperature Prediction
The NOAA OISST v2 High Resolution Dataset (Huang et al., 2021) is a long-term Climate Data Record that integrates observations from multiple platforms (satellites, ships, buoys, and Argo floats) into a global gridded product. For this experiment, we use daily averages of sea surface temperature in the Gulf of Mexico between $1981$ and $2024$, represented as $64\times 64$ single-channel images. We cluster the measurements over five-year periods and select five representative months to define five time frames: January ($t_{0}$), March ($t_{1}$), May ($t_{2}$), July ($t_{3}$), and September ($t_{4}$). Each month corresponds to a distribution of images, denoted as $\{{\rho}_{a}\ \text{(January)},{\rho}_{m_{1}}\ \text{(March)},{\rho}_{m_{2}}\ \text{(May)},{\rho}_{m_{3}}\ \text{(July)},{\rho}_{b}\ \text{(September)}\}$. A sample from one of these distributions is a heatmap of the Gulf’s temperature for a specific day in the corresponding month of the specified five-year period. The goal of this test is to evaluate our method’s ability to interpolate across missing time frames, generating realistic temperature maps for months without data. To this end, we partition the dataset into a training set $\{t_{0},t_{2},t_{4}\}$ and a validation set $\{t_{1},t_{3}\}$, and assess the quality of predictions on the held-out months. To encourage generalization beyond the training data, we introduce a potential function $U$ that penalizes deviations from the learned data manifold. Building on the approach of Song and Itti (2025), where generative models are evaluated by measuring the distance between their outputs and a geometric manifold of real images learned by a VAE, we adopt a similar strategy. Specifically, we use a state-of-the-art VAE architecture, with parameters listed in Table 22, to learn the manifold of the training images in our dataset. The potential function $U(x^{t})$, for samples $x^{t}\sim{\rho}^{t}$, is then defined as the squared distance between a bridge sample $x^{t}$ and its VAE-projected reconstruction $\tilde{x}^{t}=\text{VAE}(x^{t})$: $U(x^{t})=\lVert x^{t}-\tilde{x}^{t}\rVert^{2}$. Results for each five-year periods are presented below.
@@
In these results, the scalar Hamiltonian function $H^{t_{k}}$ in the pullback Jacobi metric, $T_{\theta}^{*}\tilde{g}_{\mathrm{J}}=\big(H^{t_{k}}-\mathcal{F}(\rho^{t})-\mathcal{B}(\rho^{t})\big)\,T_{\theta}^{*}g^{\mathcal{W}_{2}}$ is linearly varied from an initial value $H^{t_{0}}=1.36$ to a final value $H^{t_{K}}=1.0$. This law is treated as a hyperparameter of the methodology, and all ten experiments reported in the following tables were conducted under this Hamiltonian behavior. In the geodesic computation, the solution path seeks to minimize $T_{\theta}^{*}\tilde{g}_{\mathrm{J}}$. The functional $\mathcal{F}(\rho^{t})$ measures the distance with respect to the data manifold, i.e., it evaluates whether the samples from the bridge remain coherent with the images observed during training, and it is minimized along the trajectory. Since $H^{t_{k}}$ is prescribed, the entropy term $\mathcal{B}(\rho^{t})$ is effectively maximized up to the limit set by $H^{t_{k}}$, as the metric cannot become negative. Hence, the assigned Hamiltonian energy $H^{t_{k}}$ determines the admissible entropy level of the bridge. In this experiment, a decreasing trend of $H^{t_{K}}$, corresponding to an increase in physical energy (see Appendix E.4), was observed to be beneficial for modeling the warmer months, which appear to be more distant from the rest of the dataset in terms of Wasserstein distances compared to the colder months. A higher final energy $H^{t_{K}}$, and thus a higher final entropy $\mathcal{B}$, was observed to be beneficial for modeling the warmer months. This effect can be associated with the increased thermodynamic entropy of such cases, leading to more diverse samples and larger variability across the data manifold. By prescribing a higher energy level, the model is encouraged to capture this diversification, exploring regions of the data manifold that in other contexts (Arvanitidis et al., 2018) are regarded as uncertain and are typically avoided.
@@
Table 22: Architecture of the ConvVAE used in the Sea Temperature Prediction experiment. All Conv2D and ConvTranspose2D layers use kernel size $4$, stride 22, and padding 11, followed by ReLU activations (except the last layer, which uses Sigmoid).
@@
Metric CWG GSBM DSBM SB-Flow
@@
Metric CWG GSBM DSBM SB-Flow
@@
Metric CWG GSBM DSBM SB-Flow
@@
Metric CWG GSBM DSBM SB-Flow
@@
Metric CWG GSBM DSBM SB-Flow
@@
Metric CWG GSBM DSBM SB-Flow
@@
Metric CWG GSBM DSBM SB-Flow
@@
Metric CWG GSBM DSBM SB-Flow
@@
Metric CWG GSBM DSBM SB-Flow